OpenStudy (anonymous):

Question 4 (Multiple Choice Worth 4 points) [2.03] Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 8 − 5x = 4x − 1 4x + 4 = 10 − 2x The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describes solutions to equations shown.

OpenStudy (anonymous):

@waterineyes can you help?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

The solutions seem to be the same because 8-5x=4x-1 . add -5x with 4x n u get 9x so 8=9x-1. move -1 to the other side n get 9 9=9x x=1. 4x+4=10-2x. move -2x to other side to get 6x+4=10. subtract 4 from 10 and get 6 so 6x=6 x=1. hope it helps